OpenStudy (anonymous):

factor the GCF and then the resulting trinomial: 3x^3-3x^2-6x

OpenStudy (anonymous):

GCF = 3x

OpenStudy (anonymous):

3x^3-3x^2-6x = 3x(x^2 - x - 2) factoring x^2 - x -2 = x^2 -2x + x -2 = x(x - 2) +1(x - 2) = (x + 1)(x - 2) so 3x^3-3x^2-6x = 3x(x^2 - x - 2) = (3x)(x + 1)(x - 2)

OpenStudy (ash2326):

We have the trinomial \[3x^3-3x^2-6x\] We need to factor it, Let's take out x from all the terms \[x(3x^2-3x-6)\] We'll now find the factors of \[3x^2-3x-6)\] Let's take out 3 common \[3(x^2-x-2)\] we have now \[x^2-x-2]\ We need to find factors of -2 such that their sum is -1 these are -2 and 1 so \[x^2-2x+x-2\] We have \[x(x-2)+1(x-2)\] So we get \[(x-2)(x+1) So the factors of \(3x^3-3x^2-6x\) are \[3x(x-2)(x+1)\]